OpenStudy (anonymous):

does anyone take algebra 1B? Im stuck and need help with a question

OpenStudy (anonymous):

i dont entirely understand what this question means or howto solve it? 6. What is the relationship between the exponent of the base 3 and the value of n?

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Do you have some sort of pic or equation with that?

OpenStudy (anonymous):

Yes Consider the sequence 2, 6, 18, 54 .... Let n= the term number in the sequence Let A(n)= The value of the nth term of the sequence 4. What is the common ratio in the sequence? 5. Complete each statement a. A(1) = 2 = 2*3 A(1) = 2*3^1 b. A(2) = 6 = 2*3 = 2*3 A(2)= 2*3^1 c. A(3) = 18 = 2*3*3 A(3) = 2*3^2 d. A(4) = 54 = 2*3*3*3 A(4) = 2*3^3 i got the answers ^^ i just need help with that 1 question

OpenStudy (anonymous):

So, basically the power to the n: Like this: \[3^n\] is how many times that number is multiplied with itself. So 3^1 = 3 So 3^2 = 3*3 (2 times) 3^3 = 3*3*3 3^4 = 3*3*3*3 And so on, do you see it?

OpenStudy (anonymous):

Ohh no i understand it thank you!

OpenStudy (anonymous):

No problem, hope it helped, mind marking best response?

OpenStudy (anonymous):

Oh, and anynumber to the 0 power equals to 1